# Sparse Portfolio Selection via Topological Data Analysis based Clustering

### v1

@@ 3.1 Stability of Persistence diagrams under new distance measures
Proposition 3.2.
(Cohen-Steiner et al. (2007),Kusano et al. (2017),Chazal et al. (2014)) Let $(E,d_{E})$ be a metric space and $X$ and $Y$ be finite subsets of $E$. Then the persistence diagrams ${\mathcal{D}}_{X}$ and ${\mathcal{D}}_{Y}$ satisfy
@@
Proposition 3.3.
Cohen-Steiner et al. (2010),Kusano et al. (2017) Let $1\leq\hat{p}\leq p<\infty$, and ${\mathcal{D}}_{X}$ and ${\mathcal{D}}_{Y}$ be persistence diagrams whose degree $p$ total persistences [^7] Given a persistence diagram ${\mathcal{D}}_{X}=\{a_{j}=(b_{j},d_{j})\}$, the degree $p$ total persistence of ${\mathcal{D}}_{X}$ is defined as sum of persistence of all the points of degree $p$ in ${\mathcal{D}}_{X}$ i.e., $Pers_{p}({\mathcal{D}}_{X})=\sum_{j}(d_{j}-b_{j})$. are bounded from above. Then,
@@
Proposition 3.4.
Kusano et al. (2017) Let $E$ be a triangulable compact subspace in ${\mathbb{R}}^{d}$, ${X}$ be a finite subset of $E$, and $p>d$. Then,
@@
By definition, for any choice of $\eta$ and $l$, $X_{i},~{}Y_{i};~{}i\in\{1,2,\ldots,m\}$, which represents the point cloud of $i$th component in $seq({x})$ and $seq({y})$, say ${x}_{i}$ and ${y}_{i}$, respectively, are again subsets of $E$. Therefore, using Proposition 3.2, we have
@@
Using Proposition 3.3 and 3.4, we have
@@
Remark 3.2.
The $DWD_{p}$ is a special case of $WD_{p}$ with one component as zero $PD$. Therefore, the stability of persistence diagrams with respect to $DWD_{p}$ is trivial. Further, the stability of the proposed distances on persistence landscapes follows directly from the landscape stability theorems, Theorem 12 - Theorem 16 in Bubenik et al. (2015). Further, the distance metrics $AWD_{p},DWD_{p},ALD_{p}$ and $DLD_{p}$ are pseudo-metrics. The properties of non-negativity and symmetry are easily verified. Also, the triangular inequality property is satisfied since they are defined using the Wasserstein and $p$-landscape distance, which are again pseudo-metrics.
@@ 4 Our Methodology for Sparse Portfolios
where $\bm{\mu}=(\mu_{1},\ldots,\mu_{n})$ is the sample mean vector and $\bm{\Sigma}$ is the sample variance-covariance matrix of the asset returns calculated over the observation period, $t=0,1,2\ldots,T$. Further, $\bm{w}^{\prime}\bm{\mu}$ and $\bm{w}^{\prime}\bm{\Sigma}\bm{w}$ denote the return and the variance of the portfolio. The parameter $\gamma$ is the risk aversion parameter that controls the trade-off between the portfolio’s risk and return. We follow DeMiguel et al. (2009b) and set $\gamma=1$ for the empirical analysis.
@@
The most common choice for the tracking error $TE(\bm{w})$ in the literature is the mean of the squared deviations of the portfolio return and the index return, traditionally called the tracking error variance ($TEV$). We also adopt $TEV$ as the tracking error measure  [^8] One of the key benefits of employing this metric for tracking error is that the objective function is quadratic in the asset weights, allowing for using standard and very efficient quadratic programming algorithms to minimize it. A further advantage of this tracking error is that it avoids the issues associated with metrics based solely on the variance of the deviation since these measures ignore the possibility of a constant difference between the two returns (Beasley et al. (2003)). Instead of the squared deviations, the absolute error is also a popular choice in the literature (e.g., Prigent (2007); Rudolf et al. (1999)). This is advantageous because the optimization problem reduces to a linear program that can be efficiently solved. In practice, squared deviation is preferred over absolute errors as a statistical criterion, especially when large errors are undesirable. which is defined as
@@ 4.1 TDA-based Sparse Portfolios
In this paper, we employ affinity propagation clustering (APC) Redmond et al. (2011) to obtain clusters. Our choice of adopting APC mainly stems from the requirement that the representative center of the clusters must be a part of the data set (refer to Strategy 2). Further, we should get the same result, i.e., the same clusters and centers, if we run our strategy on the same data set but on different platforms/software. Clustering techniques, such as $K$-Means and spectral clustering, assign random data points as initial cluster centers and shift these centers in each iteration. The drawback of these methods is that they yield different clustering sets upon each execution due to random initialization, and the centroids often lack a meaningful connection to actual variables in the dataset Jain et al. (1999). Affinity propagation addresses this issue by directly returning exemplars or centroids that correspond to actual data points.
Moreover, APC demonstrates superior accuracy compared to other clustering algorithms (Redmond et al. (2011)) such as K-means MacQueen et al. (1967), and spectral clustering Shi and Malik (2000); Murzin et al. (1995). Notably, APC doesn’t require the pre-specification of the number of clusters, unlike k-means and spectral clustering. Consequently, the similarity matrix is the only input needed for affinity propagation clustering. Finally, as we deal with time series of stock returns, which typically behaves like white noise, clustering algorithms such as K-means would not be suitable as demonstrated in Zheng et al. (2020).
@@
where $\sigma$ is a free parameter. [^9] These are kernels in strict sense Paulsen and Raghupathi (2016) and induce similarity matrices which are semi-positive definite and hence can be used as input for the clustering algorithms. It is worth highlighting that apart from the advantage of the new TDA-based distance measure over $p$-Wasserstein distance that it allows to compare time-series data as discussed in Section 3, it offers a substantial computational speedup as it reduces the computational complexity to compute the persistence landscape for a large time series by dividing it into computations over several sub-series. Further, this division into several sub-series accounts for heteroscedasticity and non-stationarity in the financial data.
@@
where $d_{i}(x,y)=\sqrt{2(1-\rho_{i}(x,y))}~{}~{}i\in\{1,2\}$, with $\rho_{1}$ and $\rho_{2}$ denoting the Spearman’s rank correlation and Pearson’s correlation, respectively. [^10] The distance measure $d_{1}$ defined using Pearson’s correlation is the most commonly used distance measure in the literature Onnela et al. (2003). However, because these $x,y$ are logarithmic returns in our scenario, Spearman’s or Kendall’s rank correlation coefficient is a much better choice due to its robustness Zheng et al. (2020). As a result, we select both of these correlation measures for comparison. We use the same notation $K_{i}~{}i=1,\ldots,6$ to represent the kernels and the corresponding similarity matrix.
The kernels, $K_{1}$ to $K_{6}$, depend on hyper-parameter $\sigma$, which needs to be estimated from training data. This paper adopts a data-driven approach for estimating $\sigma$, specifically employing the local scaling algorithm Yang et al. (2008); Shang et al. (2012). More precisely, we use a local scaling parameter $\sigma_{ij}$ for each pair of input data $(x_{i},x_{j})$ defined as
@@
where $x_{m}(x_{\cdot})$ denotes the $m$th nearest neighbor of the input $x_{\cdot}$ in terms of distance measured by $d$ used in the corresponding kernel. [^11] The selection of $m$ is independent of scale. We take $m=7$ in our analysis since it provides good clustering results Yang et al. (2008). The adoption of diverse scaling parameters for each pair $(x_{i},x_{j})$ rather than a single scaling parameter $\sigma$ allows us to exploit the local statistics of the neighborhoods around the corresponding input pair, thereby facilitating a data-driven self-tuning of $\sigma$.
@@ 4.2 Performance Measures
It measures how closely the optimally tracking portfolio replicates the benchmark index. Further, we also apply one-tailed $t$-test, with null hypothesis $H_{0}:\,\text{TE}_{s_{1}}-\text{TE}_{s_{2}}\leq 0$ against the alternative hypothesis $H_{a}:\text{TE}_{s_{1}}-\text{TE}_{s_{2}}>0$, to test whether the TE from strategy $s_{1}$ is statistically indistinguishable to that from strategy $s_{2}$.121212 The $t$-test is used to test if the mean $\mu$ of the population is ($\geq$  or  $\leq$) to some constant $\mu_{0}$ when the variance of the population is unknown. To test the following hypothesis $H_{0}:\mu\leq\mu_{0}~{}~{}~{}\mbox{against}~{}~{}~{}H_{a}:\mu>\mu_{0},$ the test statistics is given by $t=\frac{\bar{x}-\mu_{0}}{\frac{S}{\sqrt{N}}},$ where $S$ and $\bar{x}$ denote the sample variance and mean, respectively, and $N$ is the sample size. The critical region for the test is $t>t_{\alpha,N-1},$ where $\alpha$ is the significance level, and $t_{\alpha,N-1}$ denotes the critical value of the $t$-distribution with $N-1$ degrees of freedom. The $p$-value can be obtained as $P(t_{N-1}>t)$ and we reject the null hypothesis if $p-$value is less than $\alpha$. Second, we calculate excess mean return (EMR), defined as the mean of the difference between the out-of-sample returns from the optimal tracking portfolio and the returns of the benchmark index, that is, $\text{EMR}=\frac{\sum_{t=1}^{T}(R_{t}(\bm{w})-R_{0,t})}{T}.$ A positive (negative) EMR indicates that the tracking portfolio outperforms (underperforms) the benchmark index.
@@
Similarly, given the time series of daily out-of-sample returns obtained from mean-variance and minimum variance optimal portfolios, we follow DeMiguel et al. (2009b) and compute four quantities to assess the performance of the portfolios. First, we measure the out-of-sample mean, which is given by $\mu={\sum_{t=1}^{T}\frac{R_{t}(\bm{w})}{T}}.$ To test whether the mean from the portfolio is statistically distinguishable, we also compute the p-value of the difference using the $t$-test. We apply a one-tailed $t$-test on null hypothesis $H_{0}:\,\mu_{s_{1}}-\mu_{s_{1}}=0$ against the alternative hypothesis $H_{a}:\mu_{s_{1}}-\mu_{s_{1}}>0$ to test if the out-of-sample mean return from $s_{1}$ is statistically significant from, $s_{2}$.
@@ 5.1 Data
We consider the daily closing prices of constituents of the S&P 500 index, obtained from the Thomson Reuters datastream spanning a decade from 1st Dec 2009 to 31st Dec 2019 (2631 observations). We chose the data period according to the availability of a maximum number of constituents, resulting in 451 stocks in our dataset. We utilize daily log returns to obtain an input cloud for filtration and, consequently, to compute the distance measures for clustering. At the same time, we employ simple returns to solve the portfolio optimization problem.
@@
For out-of-sample performance analysis, we follow a rolling window scheme. [^14] We look at the problem of constructing a sparse optimal portfolio for a period $1,\ldots,T$, which corresponds to the recent past (the in-sample period) and where the path of asset returns is known. The ultimate aim is to create portfolios close to optimum performance in the $T+1,\ldots,T+L$ period (the out-of-sample period). To elaborate, we begin our first window with a training period of 126 trading days (six months) as the in-sample period to create the optimal portfolio and track its performance over a testing period of 21 trading days (one month) without any re-balancing during this period. [^15] We follow the standard of 21 trading days in a month (see, for instance, Fastrich et al. (2014)). In the next window, we shift forward our training and testing period by one month to include more recent data and exclude the oldest data points such that a fixed size of the rolling window (seven months) is maintained. We then obtain an optimal portfolio on the new training period and track its performance over the new testing period. We repeat this procedure until the entire data sample is exhausted, consisting of 119 windows (test months) and thus resulting in a time series of portfolio returns, i.e., a one-month time series corresponding to each window. Finally, we analyze the performance of the optimal portfolios over the testing period, thereby offering a comprehensive assessment of the effectiveness of the filtering strategy. The rolling window procedure is also illustrated in Figure 5.
Figure 5:  Example of a rolling window scheme: In this illustration, the orange and blue hues indicate the in-sample and out-of-sample periods, respectively. For a fixed in-sample period of length $d_{1}$ and out-of-sample period of length $d_{2}$, the first in-sample period (rolling window 1) stands for $1,\ldots,d_{1}$ with corresponding out-of-sample period $d_{1}+1,\ldots,d_{1}+d_{2}$; the next in-sample period (rolling window 2) gets shifted by $d_{2}$ to obtain an in-sample window $1+d_{2}.\ldots,d_{1}+d_{2}$, with the accompanying out-of-sample period $d_{1}+d_{2}+1,\ldots,d_{1}+d_{2}+d_{2}$, and so on. For our analysis, $d_{1}=126$ and $d_{2}=21$.

### v2

@@ 3.1 Stability of Persistence diagrams under new distance measures
Proposition 1.
([20],[49],[17]) Let $(E,d_{E})$ be a metric space and $X$ and $Y$ be finite subsets of $E$. Then the persistence diagrams ${\mathcal{D}}_{X}$ and ${\mathcal{D}}_{Y}$ satisfy
@@
Proposition 2.
[21],[49] Let $1\leq\hat{p}\leq p<\infty$, and ${\mathcal{D}}_{X}$ and ${\mathcal{D}}_{Y}$ be persistence diagrams whose degree $p$ total persistences [^6] Given a persistence diagram ${\mathcal{D}}_{X}=\{a_{j}=(b_{j},d_{j})\}$, the degree $p$ total persistence of ${\mathcal{D}}_{X}$ is defined as sum of persistence of all the points of degree $p$ in ${\mathcal{D}}_{X}$ i.e., $Pers_{p}({\mathcal{D}}_{X})=\sum_{j}(d_{j}-b_{j})$. are bounded from above. Then,
@@
Proposition 3.
[49] Let $E$ be a triangulable compact subspace in ${\mathbb{R}}^{d}$, ${X}$ be a finite subset of $E$, and $p>d$. Then,
@@
By definition, for any choice of $\eta$ and $l$, $X_{i},~{}Y_{i};~{}i\in\{1,2,\ldots,m\}$, which represents the point cloud of $i$th component in $seq({x})$ and $seq({y})$, say ${x}_{i}$ and ${y}_{i}$, respectively, are again subsets of $E$. Therefore, using Proposition 1, we have
@@
Using Proposition 2 and 3, we have
@@
Remark 2.
The $DWD_{p}$ is a special case of $WD_{p}$ with one component as zero $PD$. Therefore, the stability of persistence diagrams with respect to $DWD_{p}$ is trivial. Further, the stability of the proposed distances on persistence landscapes follows directly from the landscape stability theorems, Theorem 12 - Theorem 16 in [14]. Further, the distance metrics $AWD_{p},DWD_{p},ALD_{p}$ and $DLD_{p}$ are pseudo-metrics. The properties of non-negativity and symmetry are easily verified. Also, the triangular inequality property is satisfied since they are defined using the Wasserstein and $p$-landscape distance, which are again pseudo-metrics.
@@ 4 Our Methodology for Sparse Portfolios
where $\bm{\mu}=(\mu_{1},\ldots,\mu_{n})$ is the sample mean vector and $\bm{\Sigma}$ is the sample variance-covariance matrix of the asset returns calculated over the observation period, $t=0,1,2\ldots,T$. Further, $\bm{w}^{\prime}\bm{\mu}$ and $\bm{w}^{\prime}\bm{\Sigma}\bm{w}$ denote the return and the variance of the portfolio. The parameter $\gamma$ is the risk aversion parameter that controls the trade-off between the portfolio’s risk and return. We follow [25] and set $\gamma=1$ for the empirical analysis.
@@
The most common choice for the tracking error $TE(\bm{w})$ in the literature is the mean of the squared deviations of the portfolio return and the index return, traditionally called the tracking error variance ($TEV$). We also adopt $TEV$ as the tracking error measure  [^7] One of the key benefits of employing this metric for tracking error is that the objective function is quadratic in the asset weights, allowing for using standard and very efficient quadratic programming algorithms to minimize it. A further advantage of this tracking error is that it avoids the issues associated with metrics based solely on the variance of the deviation since these measures ignore the possibility of a constant difference between the two returns ([9]). Instead of the squared deviations, the absolute error is also a popular choice in the literature (e.g., [67, 71]). This is advantageous because the optimization problem reduces to a linear program that can be efficiently solved. In practice, squared deviation is preferred over absolute errors as a statistical criterion, especially when large errors are undesirable. which is defined as
@@ 4.1 TDA-based Sparse Portfolios
In this paper, we employ affinity propagation clustering (APC) ([70]) to obtain clusters. [^8] The details on how APC works is presented in Appendix A. Our choice of adopting APC mainly stems from the requirement that the representative center of the clusters must be a part of the data set (refer to Strategy 2). Further, we should get the same result, i.e., the same clusters and centers, if we run our strategy on the same data set but on different platforms/software. Clustering techniques, such as $k$-means and spectral clustering, assign random data points as initial cluster centers and shift these centers in each iteration. The drawback of these methods is that they yield different clustering sets upon each execution due to random initialization, and the centroids often lack a meaningful connection to actual variables in the dataset ([43]). Affinity propagation addresses this issue by directly returning exemplars that correspond to actual data points.
Moreover, APC demonstrates superior accuracy compared to other clustering algorithms ([70]) such as $k$-means ([54]), and spectral clustering ([75, 61]). Notably, APC doesn’t require the pre-specification of the number of clusters, unlike $k$-means and spectral clustering. Consequently, the similarity matrix is the only input needed for affinity propagation clustering. However, it requires the choice of $\lambda$ and $P$ (refer Appendix A for details on these parameters). We chose P as the median of the similarity matrix, a common choice in the literature, as it yields a moderate number of clusters. We run the clustering algorithm for various values of $\lambda$ and select the optimal $\lambda$ based on the Silhouette score of clusters. Finally, as we deal with time series of stock returns, which typically behaves like white noise, clustering algorithms such as $k$-means would not be suitable as demonstrated in [84]. [^9] We also compare the portfolio performance from APC to that from $K-$medoids and Hierarchical clustering in the Section 5.4
@@
where $\sigma$ is a free parameter. [^10] These are kernels in strict sense [65] and induce similarity matrices which are semi-positive definite and hence can be used as input for the clustering algorithms. It is worth highlighting that apart from the advantage of the new TDA-based distance measure over $p$-Wasserstein distance that it allows to compare time-series data as discussed in Section 3, it offers a substantial computational speedup as it reduces the computational complexity to compute the persistence landscape for a large time series by dividing it into computations over several sub-series. Further, this division into several sub-series accounts for heteroscedasticity and non-stationarity in the financial data.
@@
where $d_{i}(x,y)=\sqrt{2(1-\rho_{i}(x,y))}~{}~{}i\in\{1,2\}$, with $\rho_{1}$ and $\rho_{2}$ denoting the Spearman’s rank correlation and Pearson’s correlation, respectively. [^11] The distance measure $d_{1}$ defined using Pearson’s correlation is the most commonly used distance measure in the literature [63]. However, because these $x,y$ are logarithmic returns in our scenario, Spearman’s or Kendall’s rank correlation coefficient is a much better choice due to its robustness [84]. As a result, we select both of these correlation measures for comparison. We use the same notation $K_{i}~{}i=1,\ldots,6$ to represent the kernels and the corresponding similarity matrix.
The kernels, $K_{1}$ to $K_{6}$, depend on hyper-parameter $\sigma$, which needs to be estimated from training data. This paper adopts a data-driven approach for estimating $\sigma$, specifically employing the local scaling algorithm [83, 74]. More precisely, we use a local scaling parameter $\sigma_{ij}$ for each pair of input data $(x_{i},x_{j})$ defined as
@@
where $x_{m}(x_{\cdot})$ denotes the $m$th nearest neighbor of the input $x_{\cdot}$ in terms of distance measured by $d$ used in the corresponding kernel. [^12] The selection of $m$ is independent of scale. We take $m=7$ in our analysis since it provides good clustering results [83]. The adoption of diverse scaling parameters for each pair $(x_{i},x_{j})$ rather than a single scaling parameter $\sigma$ allows us to exploit the local statistics of the neighborhoods around the corresponding input pair, thereby facilitating a data-driven self-tuning of $\sigma$.
@@ 4.2 Performance Measures
It measures how closely the optimally tracking portfolio replicates the benchmark index. Further, we also apply one-tailed $t$-test, with null hypothesis $H_{0}:\,\text{TE}_{s_{1}}-\text{TE}_{s_{2}}\leq 0$ against the alternative hypothesis $H_{a}:\text{TE}_{s_{1}}-\text{TE}_{s_{2}}>0$, to test whether the TE from strategy $s_{1}$ is statistically indistinguishable to that from strategy $s_{2}$.131313 The $t$-test is used to test if the mean $\mu$ of the population is ($\geq$  or  $\leq$) to some constant $\mu_{0}$ when the variance of the population is unknown. To test the following hypothesis $H_{0}:\mu\leq\mu_{0}~{}~{}~{}\mbox{against}~{}~{}~{}H_{a}:\mu>\mu_{0},$ the test statistics is given by $t=\frac{\bar{x}-\mu_{0}}{\frac{S}{\sqrt{N}}},$ where $S$ and $\bar{x}$ denote the sample variance and mean, respectively, and $N$ is the sample size. The critical region for the test is $t>t_{\alpha,N-1},$ where $\alpha$ is the significance level, and $t_{\alpha,N-1}$ denotes the critical value of the $t$-distribution with $N-1$ degrees of freedom. The $p$-value can be obtained as $P(t_{N-1}>t)$ and we reject the null hypothesis if $p-$value is less than $\alpha$. Second, we calculate excess mean return (EMR), defined as the mean of the difference between the out-of-sample returns from the optimal tracking portfolio and the returns of the benchmark index, that is, $\text{EMR}=\frac{\sum_{t=1}^{T}(R_{t}(\bm{w})-R_{0,t})}{T}.$ A positive (negative) EMR indicates that the tracking portfolio outperforms (underperforms) the benchmark index.
@@
Similarly, given the time series of daily out-of-sample returns obtained from mean-variance and minimum variance optimal portfolios, we follow [25] and compute four quantities to assess the performance of the portfolios. First, we measure the out-of-sample mean, which is given by $\mu={\sum_{t=1}^{T}\frac{R_{t}(\bm{w})}{T}}.$ To test whether the mean from the portfolio is statistically distinguishable, we also compute the p-value of the difference using the $t$-test. We apply a one-tailed $t$-test on null hypothesis $H_{0}:\,\mu_{s_{1}}-\mu_{s_{1}}=0$ against the alternative hypothesis $H_{a}:\mu_{s_{1}}-\mu_{s_{1}}>0$ to test if the out-of-sample mean return from $s_{1}$ is statistically significant from, $s_{2}$.
@@ 5.1 Data
We consider the daily closing prices of constituents of the S&P 500 index, obtained from the Thomson Reuters datastream spanning a period from 1st Dec 2009 to 31st August 2022.151515We chose the constituents of the S&P 500 index as it includes the largest and the most liquid companies in the US markets and hence is widely used as a benchmark by both academics and practitioners alike. Further, the S&P 500 index is relatively stable and has least distortions driven by idiosyncrasies of small and less liquid stocks, whose behavior is not a representative of market-wide behavior. Thus, it provides a reliable dataset over a long duration that has minimum distortion due to less frequent re-balancing. Moreover, it provides a robust proxy to analyze economic trends and broader dynamics of the markets due to its comprehensive coverage of large as well as well-established companies spanning various sectors. At last, institutional investors closely monitor the performance of S&P 500 constituents, thus, making the index a key indicator of market sentiments and risk perceptions, particularly during major financial events. For completeness, we have also tested the performance of our TDA-based sparse index tracking portfolios on the benchmark data sets from [9] and [15] for index tracking problems (refer Table 8). These data sets are drawn from several major world equity indices, and thus, analyzing the performance over these data sets gives additional evidence on the adequacy and effectiveness of the TDA-based portfolios. The selection of the period stems from various reasons. First, we chose the data period according to the availability of a maximum number of constituents, resulting in 451 stocks in our dataset. Second, it covers a range of particularly significant market conditions, starting from the recovery after the sub-prime crisis to the onset of unprecedented COVID-19 pandemic, when the global markets experienced an unmatched turmoil, followed by a historic rebound to pre-crisis peaks and a stable period afterwards. Thus, the selected period allows to assess the performance of TDA across a wide range of economic regimes, including both normal and highly volatile scenarios. The COVID-19 crisis period, in particular, provides a perfect environment to test the robustness of our approach under extreme volatility.
We divide the study period into three non-overlapping periods. The first period covers a relatively stable period starting in December 2009 to July 2019. The second period, August 2019 to August 2020, allows to evaluate the robustness of TDA based approach during a crucial time of COVID-19 pandemic. The remaining period, September 2020 to August 2022 is utilized for comparison purposes, in particular, to other clustering approaches. There are several reasons for selecting August 2019 to August 2020 for COVID-19 analysis. First, choosing August 2019 as the starting month allows us to use February 2020 as the first test month. Second, it covers an extraordinary financial market event, where markets across the globe faced an unmatched turmoil, resulting in a sharp decline of 34% of the S&P500 from its pre-crisis peak in February 2020. This is in contrast to a rather slow paced decline during the sub-prime crisis. Third, as pointed out in [1], markets encountered a historic rebound to pre-crisis levels by mid-August 2020, owing to the policy responses such as liquidity support, cuts in the interest rates, emergency credit availability, thus, showing one of the fastest recoveries. Therefore, the selected period allows us to assess the models’ performance across various phases of the market, including extreme volatility and sharp recovery, thus offering insights into models’ performance under stressed scenarios. We utilize daily log returns to obtain an input cloud for filtration and, consequently, to compute the distance measures for clustering. At the same time, we employ simple returns to solve the portfolio optimization problem.
@@
For out-of-sample performance analysis, we follow a rolling window scheme. [^16] We look at the problem of constructing a sparse optimal portfolio for a period $1,\ldots,T$, which corresponds to the recent past (the in-sample period) and where the path of asset returns is known. The ultimate aim is to create portfolios close to optimum performance in the $T+1,\ldots,T+L$ period (the out-of-sample period). To elaborate, we begin our first window with a training period of 126 trading days (six months) as the in-sample period to create the optimal portfolio and track its performance over a testing period of 21 trading days (one month) without any re-balancing during this period. [^17] We follow the standard of 21 trading days in a month (see, for instance, [30]). In the next window, we shift forward our training and testing period by one month to include more recent data and exclude the oldest data points such that a fixed size of the rolling window (seven months) is maintained. We then obtain an optimal portfolio on the new training period and track its performance over the new testing period. We repeat this procedure until the entire data sample is exhausted. Finally, we analyze the performance of the optimal portfolios over the testing period, thereby offering a comprehensive assessment of the effectiveness of the filtering strategy.